A Flat bar is loaded with boxes. Box M = 50 kg, what must be the magnitude of mass m so that the bar becomes level? Neglect the weight of the bar. What torque about each end of the bar?
1
Expert's answer
2020-11-11T08:56:18-0500

If the length of the bar is L, and its axis of rotation is at distance d from the point where mass M is attached, the value of mass m in order to balance M can be found from the equilibrium of torques:


"Mgd=mg(L-d),\\\space\\
m=M\frac{d}{L-d}."

The torque about each end of the bar is


"\tau_M=Md,\\
\tau_m=m(L-d)."